4- When the Fed conducts open-market sales,

a. it sells Treasury securities, which decreases the money supply.

b. it lends money to member banks, which decreases the money supply.

c. it borrows from member banks, which increases the money supply.

d. it sells Treasury securities, which increases the money supply.

5-  When the government levies a $100 million tax on people's income and puts the $100 million back into the economy in the form of a spending program such as new interstate highway construction, the

a. tax, then, generates a $100 million decline in national income

b. tax multiplier overpowers the income multiplier, triggering a rollback in national income

c. effect on national income is uncertain

d. level of national income expands by $100 million

6- Suppose that banks desire to hold no excess reserves. If the reserve requirement is 5 percent and a bank receives a new deposit of $400, it

a. must increase required reserves by $20.

b. will initially see reserves increase by $400.

c. will be able to use this deposit to make new loans amounting to $380.

d. All of the above are correct.
Answer #1

4) a

When the Fed conducts open-market sales, it sells Treasury securities, which decreases the money supply.

5) d

When government spending and taxes increase by same amount, multiplier is 1. This means national income increases by the amount of government spending or taxes.

6) d

Required reserves= 0.05*400= 20

Loans= 400-20= 380